A 6.50 3 102-kg elevator starts from rest and moves upward for 3.00 s with constant acceleration until it reaches its cruis- ing speed, 1.75 m/s. (a) What is the average power of the elevator motor during this period? (b) How does this amount of power compare with its power during an upward trip with constant speed
1
Expert's answer
2020-03-10T11:17:32-0400

Data:

"m = 6.5 \times 10^{2} kg"

"t=3s"

"\nu = 1.75m/s"


The acceleration of elevator at first time is:

"a=\frac{\nu - \nu0}{t}"

Motor force is:

"F = m\times(a+g) = m\times(\frac{\nu - \nu0}{t} + g)"


when the elevator moves upward with constant speed, "a = 0"

So force of motor is equal:

"Fconst = m \times g"

Pressure on its power is:

"P = Fconst \times\nu"


"F = 6.5 \times10^{2} \times (\frac{1.75 - 0}{3} + 9.8)
= 650 \times (0.583 + 9.8) = 6749N"

"Fconst = 6.5 \times10^2\times9.8=6370N"

"P = 6370 \times1.75 = 11147.5W"


Answer:

1) 6749N

2) 11147.5W